GAS PHASE STRUCTURE OF TRIFLUOROACETYL PEROXYNITRATE.

ANGELIKA HERMANN, JAN NIEMEYER, HANS-GEORG MACK, \underline{DINES CHRISTEN}, and HEINZ OBERHAMMER, Institute of Physical and Theoretical Chemistry, University of Tübingen, Germany; HELGE WILLNER, Institute of Synthetic Chemistry, University of Duisburg, Germany; MARTIN SCHÄFER and ALFRED BAUDER, Laboratory of Physical Chemistry, ETH, Zürich, Switzerland.

The molecular structure and conformational properties of trifluoroacetyl peroxynitrate, FPAN, CF_{3}C(O)OONO_{2}, were investigated in the gas phase by electron diffraction, microwave spectroscopy and quantum chemical methods. All experimental and theoretical methods show the syn conformer (C=O bond of the acetyl group syn to O-O bond) to be strongly predominant relative to the anti conformer. The O-N bond is extremely long (1.526 (10) Å), which correlates with the low bond energy and the easy formation of CF_{3}C(O)OO and NO_{2} radicals in the atmosphere. The O-O bond (1.408 (8) Å) is shorter than that in hydrogen peroxide (1.464 Å) and the C-O-O-N dihedral angle is close to 85^{o}.